# Learning on Model Weights using Tree Experts

### v2

@@ Abstract
The increasing availability of public models begs the question: can we train neural networks that use other networks as input? Such models allow us to study different aspects of a given neural network, for example, determining the categories in a model’s training dataset. However, machine learning on model weights is challenging as they often exhibit significant variation unrelated to the models’ semantic properties (nuisance variation). Here, we identify a key property of real-world models: most public models belong to a small set of Model Trees, where all models within a tree are fine-tuned from a common ancestor (e.g., a foundation model). Importantly, we find that within each tree there is less nuisance variation between models. Concretely, while learning across Model Trees requires complex architectures, even a linear classifier trained on a single model layer often works within trees. While effective, these linear classifiers are computationally expensive, especially when dealing with larger models that have many parameters. To address this, we introduce Probing Experts (ProbeX), a theoretically motivated and lightweight method. Notably, ProbeX is the first probing method specifically designed to learn from the weights of a single hidden model layer. We demonstrate the effectiveness of ProbeX by predicting the categories in a model’s training dataset based only on its weights. Excitingly, ProbeX can also map the weights of Stable Diffusion into a shared weight-language embedding space, enabling zero-shot model classification.
@@ 1 Introduction
In recent years, the number of publicly available neural network models has skyrocketed, with over one million models now hosted on Hugging Face. This growth raises a new research question: can we develop neural networks capable of using other networks as input data? The emerging field of weight-space learning [31, 25, 41] studies how to design and train metanetworks, neural networks that take the weights of other neural networks as inputs (see Fig. 2). Previous works learned metanetworks that predict training data attributes [34, 24], model performance [25], and even generate new model parameters [6, 27]. In this work, we focus on predicting the categories in a model’s training dataset, e.g., the classes it was trained on. Notably, in some cases, we can embed model weights into a shared space with language, enabling tasks like zero-shot model classification.
Figure 1: Growth in Hugging Face models: The explosive growth of public models makes model weights a new data modality.
Figure 2: Weight-space learning: Left. Model weights are a direct product of the optimization process and training data. Center. Models are uploaded to public repositories, e.g., Hugging Face, often lacking details on training data or hyper-parameters. Right. Weight-space learning treats each model as a data point and designs metanetworks: networks that process weights of other models as inputs. We train metanetworks that predict categories in a model’s training dataset. We also align weights with language, enabling zero-shot model classification.
However, extracting meaningful information from model weights is challenging. While the weights of a neural network are a function of its training data, they are also affected by the optimization process, which may introduce nuisance variation unrelated to attributes of interest. Neuron permutation [14] is perhaps the most studied nuisance factor and has driven research into permutation-invariant architectures [24, 26, 25, 22, 41] and carefully designed data augmentations [34, 31, 32, 15]. In this paper, we highlight another important nuisance factor, the weights at the beginning of optimization.
@@
To evaluate our method, we introduce a dataset of $14{,}000$ models across $5$ disjoint Model Trees spanning multiple architectures and functionalities. ProbeX achieves state-of-the-art results on the task of training category prediction, accurately identifying the specific classes within a model’s training dataset. Excitingly, ProbeX can also align fine-tuned Stable Diffusion weights with language representations. This capability enables a new task: zero-shot model classification, where models are classified via a text prompt describing their training data. On this task, ProbeX achieves an impressive $89.8\%$ accuracy.
@@ 2 Related works
Despite the popularity of neural networks, little research has explored using their weights as inputs for machine learning methods. Unterthiner et al. [36], Eilertsen et al. [8] were among the first to systematically analyze model weights to predict undocumented properties like the training dataset or generalization error. Some works aim to learn general representations [34, 31, 32, 15] for multiple properties, while others incorporate specific priors [24, 26, 25, 22, 41] to directly predict the property. A major challenge with model weights is the presence of many parameter space symmetries [14]. For instance, permuting neurons in hidden layers of an MLP doesn’t change the network output. Thus, neural networks designed to take weights as inputs need to account for these symmetries. To avoid the issue of weight symmetries, recent methods [21, 22, 15] propose using probing. In this approach, a set of probes are optimized to serve as inputs to the model and the outputs act as the model representation. However, until now, this was limited to passing the probes through the entire model and did not apply to single layers.
Other applications include generating weights [11, 1, 27, 9, 6], predicting dataset size [30], merging models [38, 35], and recovering the weights of unpublished models [17, 2].
@@ 6 Experiments
Baselines. Most state-of-the-art methods do not scale to large models with hundreds of millions of parameters. We therefore compare to the following baselines: i) StatNN [36]. This permutation-invariant baseline extracts $7$ simple statistics (mean, variance, and $5$ different quantiles) for the weights and biases of each layer. It then trains a gradient-boosted tree on the concatenated statistics. ii) Dense Expert. Training a single linear layer on the flattened raw weights. Note that this baseline produces impractically large classifiers. E.g., a single layer classifier trained to classify $SD_{1k}$ typically has $1.4B$ parameters, twice the size of the entire SD model. We also attempted to run Neural Graphs [22], but it struggled to scale to the ViT architecture even when adapted to the single-layer case (see App. H). Since all attempts yielded near-random results, we did not report this baseline.
@@ 9 Conclusion
In this paper, we identify that learning from models within a Model Tree is much simpler than learning across trees. This setting is practical as most public models belong to a few large Model Trees. We therefore introduce Probing Expert (ProbeX), a theoretically grounded architecture that scales weight-space learning to large models. As public repositories consist of multiple trees, we propose a Mixture-of-Experts approach. To validate our approach, we introduce a dataset of $14,000$ fine-tuned models. We also use ProbeX to embed model weights in a shared space with language representations, enabling text-guided zero-shot model classification.

### v3

@@ Abstract
The number of publicly available models is rapidly increasing, yet most remain undocumented. Users looking for suitable models for their tasks must first determine what each model does. Training machine learning models to infer missing documentation directly from model weights is challenging, as these weights often contain significant variation unrelated to model functionality (denoted nuisance). Here, we identify a key property of real-world models: most public models belong to a small set of Model Trees, where all models within a tree are fine-tuned from a common ancestor (e.g., a foundation model). Importantly, we find that within each tree there is less nuisance variation between models. Concretely, while learning across Model Trees requires complex architectures, even a linear classifier trained on a single model layer often works within trees. While effective, these linear classifiers are computationally expensive, especially when dealing with larger models that have many parameters. To address this, we introduce Probing Experts (ProbeX), a theoretically motivated and lightweight method. Notably, ProbeX is the first probing method specifically designed to learn from the weights of a single hidden model layer. We demonstrate the effectiveness of ProbeX by predicting the categories in a model’s training dataset based only on its weights. Excitingly, ProbeX can map the weights of Stable Diffusion into a weight-language embedding space, enabling model search via text, i.e., zero-shot model classification.
@@ 1 Introduction
In recent years, the number of publicly available neural network models has skyrocketed, with over one million models now hosted on Hugging Face. Ideally, this abundance would allow users to simply download the most suitable model for their task, thereby saving resources, reducing training time, and potentially improving accuracy. However, the lack of adequate documentation for most models makes it challenging for users to determine a model’s suitability for specific tasks. This motivates developing machine learning methods that can infer model functionality and missing documentation directly from model weights. The emerging field of weight-space learning [40, 31, 54, 52, 34] studies how to design and train metanetworks, neural networks that take the weights of other neural networks as inputs (see Fig. 2). Previous works learned metanetworks that predict training data attributes [43, 29], model performance [31], and even generate new model parameters [9, 35]. In this work, we focus on predicting the categories in a model’s training dataset as a proxy for the concepts it can recognize or generate. As a initial step towards model search-by-text, we demonstrate that it is sometimes possible to align weights with language, enabling zero-shot model classification.
Figure 1: Growth in Hugging Face models: The number of public models is growing fast, but they are mostly undocumented. Fully benefiting from them requires effective model search.
Figure 2: Weight-space learning: Left. Model weights are a direct product of the optimization process and training data. Center. Models are uploaded to public repositories, e.g., Hugging Face, typically lacking documentation. This prevents model search, making models hard to discover and reuse. Right. Weight-space learning treats each model as a data point and designs metanetworks: networks that process weights of other models as inputs. We train metanetworks that predict categories in a model’s training dataset. As a first step towards model search-by-text, we also align weights with language, enabling zero-shot model classification.
However, extracting meaningful information from model weights is challenging. While the weights of a neural network are a function of its training data, they are also affected by the optimization process, which may introduce nuisance variation unrelated to attributes of interest. Neuron permutation [17] is perhaps the most studied nuisance factor and has driven research into permutation-invariant architectures [29, 32, 31, 27, 54] and carefully designed augmentations [43, 40, 41, 18]. In this paper, we highlight another important nuisance factor, the weights at the beginning of optimization.
@@
To evaluate our method, we introduce a dataset of $14{,}000$ models across $5$ disjoint Model Trees spanning multiple architectures and functionalities. ProbeX achieves state-of-the-art results on the task of training category prediction, accurately identifying the specific classes within a model’s training dataset. Excitingly, ProbeX can also align fine-tuned Stable Diffusion weights with language representations. This capability enables a new task: zero-shot model classification, where models are classified via a text prompt describing their training data, ProbeX achieves $89.8\%$ accuracy on this.
@@ 2 Related works
Despite the popularity of neural networks, little research has explored using their weights as inputs for machine learning methods. Unterthiner et al. [47], Eilertsen et al. [11] were among the first to systematically analyze model weights to predict undocumented properties like the training dataset or generalization error. Some works aim to learn general representations [43, 40, 41, 18] for multiple properties, while others incorporate specific priors [29, 32, 31, 27, 54, 39] to directly predict the property. A major challenge with model weights is the presence of many parameter space symmetries [17]. For instance, permuting neurons in hidden layers of an MLP doesn’t change the network output. Thus, neural networks designed to take weights as inputs need to account for these symmetries. To avoid the issue of weight symmetries, recent methods [26, 27, 18, 5, 30] propose using probing. In this approach, a set of probes are optimized to serve as inputs to the model and the outputs act as the model representation. Probes are also used in other fields such as [23, 46, 2, 45, 8] mechanistic interpretability for solving different tasks. However, until now, this was limited to passing the probes through the entire model and did not apply to single layers. Concurrent to our work, Putterman et al. [36] propose a method for linearly classifying LoRA [22] models to infer their functionality. While they do not describe it this way, it can be viewed as probing.
Other applications include generating weights [14, 1, 35, 12, 9], merging models [50, 44, 25, 49, 33], and recovering the weights of unpublished models [20, 3].
@@ 6 Experiments
Baselines. Most state-of-the-art methods do not scale to large models with hundreds of millions of parameters. We therefore compare to the following baselines: i) StatNN [47]. This permutation-invariant baseline extracts $7$ simple statistics (mean, variance, and $5$ different quantiles) for the weights and biases of each layer. It then trains a gradient-boosted tree on the concatenated statistics. ii) Dense Expert. Training a single linear layer on the flattened raw weights. Note that this baseline produces impractically large classifiers. E.g., a single layer classifier trained to classify $SD_{1k}$ typically has $1.4B$ parameters, twice the size of the entire SD model. We also attempted to run Neural Graphs [27], but it struggled to scale to the ViT architecture even when adapted to the single-layer case (see App. H). Since all attempts yielded near-random results, we did not report it.
@@ 9 Conclusion
In this paper, we take the first step toward model search based solely on model weights. We identify that learning from models within the same Model Tree is significantly simpler than learning across different trees. This setting is practical as most public models belong to a few large Model Trees. We therefore introduce Probing Expert (ProbeX), a theoretically grounded architecture that scales weight-space learning to large models. As public repositories consist of multiple trees, we propose a Mixture-of-Experts approach. We demonstrate that ProbeX can embed model weights into a shared representation space alongside language embeddings, enabling text-guided zero-shot model classification.